# Toward Global Convergence of Gradient EM for Over-Parameterized Gaussian Mixture Models

### v1

@@ 1.1 Gaussian Mixture Model (GMM)
We consider the canonical Gaussian Mixture Models with weights $\bm{\pi}=(\pi_{1},\ldots,\pi_{n})$ ($\sum_{i=1}^{n}\pi_{i}=1$), means $\bm{\mu}=(\mu_{1}^{\top},\ldots,\mu_{n}^{\top})^{\top}$ and unit covariance matrices $I_{d}$ in $d$-dimensional space. Following a widely-studied setting [Balakrishnan et al., 2014, Yan et al., , Daskalakis et al., 2017], we set the weights $\bm{\pi}$ and covariances $I_{d}$ in student GMM as fixed, and the means $\bm{\mu}=(\mu_{1}^{\top},\ldots,\mu_{n}^{\top})^{\top}$ as trainable parameters. We use $\text{GMM}(\bm{\mu})$ to denote the GMM model parameterized by $\bm{\mu}$, which can be described with probability density function (PDF) $p_{\bm{\mu}}:\mathbf{R}^{d}\to\mathbf{R}_{\geq 0}$ as
@@ 1.2 Gradient EM algorithm
The focus of this paper is the non-convex optimization analysis, so we consider using population gradient EM algorithm to learn GMM (1), where the observed variable is $x\in\mathbf{R}^{d}$ and latent variable is the index of membership Gaussian in GMM. We follow the standard teacher-student setting where a student model $\text{GMM}(\bm{\mu})$ with $n\geq 2$ Gaussian components learns from data generated from a ground truth teacher model $\text{GMM}(\bm{\mu}^{*})$. We consider the over-parameterized setting where the ground truth model $\text{GMM}(\bm{\mu}^{*})$ is a single Gaussian distribution $\mathcal{N}(0,I_{d})$, namely $\bm{\mu}^{*}=({\mu_{1}^{*}}^{\top},\ldots,{\mu_{n}^{*}}^{\top})^{\top}=({0}^{\top},\ldots,{0}^{\top})^{\top}$. Our problem could be seen as a strict generalization of Dwivedi et al. [2018b], where they studied using mixture model of two Gaussians with symmetric means (they set constraint $\mu_{2}=-\mu_{1}$) to learn one single Gaussian.
@@
    $\mu_{i}(t+1)=\mu_{i}(t)-\eta\nabla_{\mu_{i}}Q(\bm{\mu}(t)|\bm{\mu}(t))=\mu_{i}(t)-\eta\mathbf{E}_{x\sim\mathcal{N}(0,I_{d})}\left[\-\psi_{i}(x|\bm{\mu}(t))(\mu_{i}(t)-x)\right].$ (3)
@@ 1.3 Loss function of gradient EM
For any $\bm{\mu}$, $\nabla Q(\bm{\mu}|\bm{\mu})=\nabla\mathcal{L}(\bm{\mu})$.
Fact 1 states that the gradient of $Q$ function that gradient EM optimizes in each iteration is identical to the gradient of loss function $\mathcal{L}$. This observation is very useful since it implies that gradient EM is equivalent to gradient descent (GD) algorithm on $\mathcal{L}$. This observation is not a new discovery of ours but actually a wide-spread folklore (see [Jin et al., 2016]). However, our new contribution is to observe Fact 1 is very helpful for analyzing gradient EM, and to construct a new convergence analysis framework for gradient EM based on it.
1.4 Notations
In this paper, we adopt the following notational conventions. We denote $\{1,2,\ldots,n\}$ with $[n]$. $\bm{\mu}=(\mu_{1}^{\top},\ldots,\mu_{n}^{\top})^{\top}\in\mathbf{R}^{nd}$ denotes the parameter vector of GMM obtained by concatenating Gaussian mean vectors $\mu_{1},\ldots,\mu_{n}$ together. For any vector $\mu$, $\mu(t)$ denotes its value at time step $t$, sometimes we omit this iteration number $t$ when its choice is clear and simply abbreviate $\mu(t)$ as $\mu$. We define a shorthand of expectation taken over the ground truth GMM $\mathbf{E}_{x\sim\mathcal{N}(0,I_{d})}[\cdot]$ as $\mathbf{E}_{x}[\cdot]$. For any vector $v\neq 0$, we use $\overline{v}\coloneqq v/\|v\|$ to denote the normalization of $v$. We define (with a slight abuse of notation) ${i_{\max}}\coloneqq\arg\max_{i\in[n]}\{\|\mu_{i}\|\}$ as the index of $\mu_{i}$ with the maximum norm, and $\mu_{\max}\coloneqq\|\mu_{{i_{\max}}}\|=\max_{i\in[n]}\{\|\mu_{i}\|\}$ as the maximum norm of $\mu_{i}$. In particular, $\mu_{\max}(t)=\max\{\|\mu_{1}(t)\|,\ldots,\|\mu_{n}(t)\|\}$. Similarly, $\pi_{\min}\coloneqq\min_{i\in[n]}\pi_{i}$ and ${\pi_{\max}}\coloneqq\max_{i\in[n]}\pi_{i}$ denotes the minimal and maximal $\pi_{i}$, respectively. We use $\nabla_{\mu_{i}}\mathcal{L}$ to denote the gradient of $\mu_{i}$ on $\mathcal{L}$, and $\nabla\mathcal{L}=(\nabla_{\mu_{1}}\mathcal{L}^{\top},\ldots,\nabla_{\mu_{n}}\mathcal{L})^{\top}$ denotes the collection of all gradients.
@@ 1.5 Technical overview
New likelihood-based analysis framework. The traditional convergence analysis for EM/gradient EM in previous works Balakrishnan et al. [2014], Yan et al. , Kwon and Caramanis [2020] proceeds by showing the distance between the model and the ground truth GMM in the parameter space contracts linearly in every iteration. This type of approach meets new challenges in the over-parameterized $n$-Gaussian mixture setting since the convergence is both sub-linear and non-monotonic. To address these problems, we propose a new likelihood-based convergence analysis framework: instead of proving the convergence of parameters, our analysis proceeds by showing the likelihood loss function $\mathcal{L}$ converges to 00. The new analysis framework is more flexible and allows us to overcome the aforementioned technical barriers.
@@ 2.2 N-Gaussian mixtures
A series of paper Yan et al. , Zhao et al. , Kwon and Caramanis [2020], Segol and Nadler follow the framework proposed by Balakrishnan et al. [2014] to prove the local convergence of EM for $n$-GMM. While their result applies to the more general $n$-Gaussian mixture ground truth setting, their framework only provides local convergence guarantee and cannot be directly applied to our setting.
@@ 3.1 Global convergence of gradient EM
Consider training a student $n$-component GMM initialized from $\bm{\mu}(0)=(\mu_{1}(0)^{\top},\ldots,\mu_{n}(0)^{\top})^{\top}$ to learn a single-component ground truth GMM $\mathcal{N}(0,I_{d})$ with population gradient EM. If the step size satisfies $\eta\leq O\left(\frac{\exp\left(-8n\mu_{\max}^{2}(0)\right)\pi_{\min}^{2}}{n^{2}d^{2}(\frac{1}{\mu_{\max}(0)}+\mu_{\max}(0))^{2}}\right)$, then gradient EM converges globally with rate
@@
where constant $\gamma=\Omega\left(\frac{\eta\exp\left(-16n\mu_{\max}^{2}(0)\right)\pi_{\min}^{4}}{n^{2}d^{2}(1+\mu_{\max}(0){\sqrt{dn}})^{4}}\right)\in\mathbf{R}^{+}$, $\mu_{\max}(0)=\max\{\|\mu_{1}(0)\|,\ldots,\|\mu_{n}(0)\|\}$.
@@
The convergence rate in Theorem 2 has a factor exponentially small in the initialization scale ($\gamma\propto\exp(-16n\mu_{\max}^{2}(0))$). We would like to stress that this is again due to algorithmic nature of the problem rather than the limitation of analysis. In Section 3.2, we prove that there exists bad regions with exponentially small gradients so that when initialized from such region, gradient EM gets trapped locally for $\exp(\Omega(\mu_{\max}^{2}(0)))$ number of steps. Therefore, a convergence speed guarantee exponentially small in square of initialization scale is inevitable and cannot be improved.
@@
Theorem 2 is fundamentally different from convergence analysis for EM/gradient EM in previous works Yan et al. , Dwivedi et al. [2019], Balakrishnan et al. [2014] which proved monotonic linear contraction of parameter distance $\|\bm{\mu}(t)-\bm{\mu}^{*}\|$. But our result also implies global convergence since loss function $\mathcal{L}$ converging to 00 is equivalent to convergence of gradient EM to MLE.
@@ 3.2 Necessity of exponentially small factor in convergence rate
In this section we prove that a factor of $\exp(-\Theta(\mu_{\max}^{2}(0)))$ is inevitable in the global convergence rate guarantee of gradient EM. To demonstrate this, we show that there exists some bad region such that initializing from this region will trap gradient EM for an exponentially long time before it converges to the global minimum. Our result is the following theorem.
@@
For any $n=2l+1$, consider gradient EM initialized at point $\mu_{1}(0)=0,\mu_{2}(0)=\cdots=\mu_{l+1}(0)=12\sqrt{d}e_{1},\mu_{l+2}(0)=\cdots=\mu_{2l+1}(0)=-12\sqrt{d}e_{1}$, where $e_{1}=(1,0,\ldots,0)^{\top}$ is a standard unit vector. Then population gradient EM will be trapped in a bad local region around $\bm{\mu}(0)$ for exponentially many number of time steps $T=\frac{1}{15n\eta}e^{d}=\frac{1}{15n\eta}\exp(\Theta(\mu_{\max}^{2}(0)))$. More rigorously, for any $0\leq t\leq T$, we have
$\|\mu_{i}(t)\|\geq 10\sqrt{d},\forall i\neq 1.$
Theorem 7 states that, when initialized from some bad points $\bm{\mu}(0)$, after $\exp(\Theta(\mu_{\max}^{2}(0)))$ number of time steps, gradient EM will still stay in this local region and remain $10\sqrt{d}$ distance away from the global minimum $\bm{\mu}=0$. Therefore an exponentially small factor in convergence rate is inevitable.
@@ 4.1 Difficulties of a global convergence proof and our new analysis framework
Proving the global convergence of gradient EM for general $n$-Gaussian mixture is highly nontrivial. While there have been many previous works [Balakrishnan et al., 2014, Yan et al., , Dwivedi et al., 2018b] studying either local convergence or the special case of $2$-Gaussian mixtures, they all focus on showing the contraction of parametric error. Namely, their proof proceeds by showing the distance between the model parameter and the ground truth contracts, usually by a fixed linear ratio, in each iteration of the algorithm. However, this kind of approach faces various challenges for our general problem where the convergence is both sublinear and non-monotonic. Since the convergence rate is sublinear (see Remark 3), showing a linear contraction per iteration is no longer possible. Since the convergence is non-monotonic [^1] To see this, consider $n=2,\mu_{1}=0,\mu_{2}=(1,0,\ldots,0)^{\top}$, then the norm of $\mu_{1}$ strictly increases after one iteration., we also cannot show a strictly decreasing parametric distance.
@@ 4.2 Proof ideas for Theorem 2
Recall that $\bar{x}=\frac{x}{\|x\|}$. See detailed derivation in (23). Using (5), combined with the properties of $\nabla_{x}\tilde{\bm{\psi}}_{\bm{\mu}}(x)$, we can obtain the following lemma.
@@
$\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]\geq\frac{\exp\left(-8\mu_{\max}^{2}\right)}{40000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}.$
@@
$\left\langle\nabla_{\bm{\mu}}Q(\bm{\mu}|\bm{\mu}),\bm{\mu}\right\rangle=\mathbf{E}_{x}[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}]=\Omega\left(\frac{\exp\left(-8\mu_{\max}^{2}\right)\pi_{\min}^{2}}{d(1+\mu_{\max}{\sqrt{d}})^{2}}\mu_{\max}^{4}\right).$
@@
The second problem is that our local smoothness theorem requires $\bm{\mu}$ to be bounded, therefore we need to show a regularity condition that for each $i$, $\bm{\mu}_{i}(t)$ stays in a bounded region during gradient EM updates. This is not easy to prove for each individual $\bm{\mu}_{i}$ due to the same non-monotonic issue mentioned in Section 4.1. To establish such a regularity condition, we introduce the following potential function.
Definition 15.
Define potential function $U:\mathbf{R}^{nd}\to\mathbf{R}\;$ for $\text{GMM}(\bm{\mu})$ as
$U(\bm{\mu})=\sum_{i\in[n]}\|\mu_{i}\|^{2}.$
We prove that the potential function $U$ remains bounded, implying each $\bm{\mu}_{i}$ remains well-behaved. With this regularity condition, combined with the previous two steps, we finish the proof of Theorem 2 via mathematical induction.

### v2

@@ 1.1 Gaussian Mixture Model (GMM)
We consider the canonical Gaussian Mixture Models with weights $\bm{\pi}=(\pi_{1},\ldots,\pi_{n})$ ($\sum_{i=1}^{n}\pi_{i}=1$), means $\bm{\mu}=(\mu_{1}^{\top},\ldots,\mu_{n}^{\top})^{\top}$ and unit covariance matrices $I_{d}$ in $d$-dimensional space. Following a widely-studied setting [Balakrishnan et al., 2014, Yan et al., 2017, Daskalakis et al., 2017], we set the weights $\bm{\pi}$ and covariances $I_{d}$ in student GMM as fixed, and the means $\bm{\mu}=(\mu_{1}^{\top},\ldots,\mu_{n}^{\top})^{\top}$ as trainable parameters. We use $\text{GMM}(\bm{\mu})$ to denote the GMM model parameterized by $\bm{\mu}$, which can be described with probability density function (PDF) $p_{\bm{\mu}}:\mathbf{R}^{d}\to\mathbf{R}_{\geq 0}$ as
@@ 1.2 Gradient EM algorithm
The focus of this paper is the non-convex optimization analysis, so we consider using population gradient EM algorithm to learn GMM (1), where the observed variable is $x\in\mathbf{R}^{d}$ and latent variable is the index of membership Gaussian in GMM. We follow the standard teacher-student setting where a student model $\text{GMM}(\bm{\mu})$ with $n\geq 2$ Gaussian components learns from data generated from a ground truth teacher model $\text{GMM}(\bm{\mu}^{*})$. We consider the over-parameterized setting where the ground truth model $\text{GMM}(\bm{\mu}^{*})$ is a single Gaussian distribution $\mathcal{N}(\mu^{*},I_{d})$, namely $\bm{\mu}^{*}=({\mu^{*}}^{\top},\ldots,{\mu^{*}}^{\top})^{\top}$. We can then further assume $w.l.o.g.$ that $\mu^{*}=0$. Our problem could be seen as a strict generalization of Dwivedi et al. [2018b], where they studied using mixture model of two Gaussians with symmetric means (they set constraint $\mu_{2}=-\mu_{1}$) to learn one single Gaussian.
@@
    $\mu_{i}(t+1)=\mu_{i}(t)+\eta\nabla_{\mu_{i}}Q(\bm{\mu}(t)|\bm{\mu}(t))=\mu_{i}(t)-\eta\mathbf{E}_{x\sim\mathcal{N}(0,I_{d})}\left[\-\psi_{i}(x|\bm{\mu}(t))(\mu_{i}(t)-x)\right].$ (3)
@@ 1.3 Loss function of gradient EM
For any $\bm{\mu}$, $\nabla Q(\bm{\mu}|\bm{\mu})=-\nabla\mathcal{L}(\bm{\mu})$.
Fact 1 states that the gradient of $Q$ function that gradient EM optimizes in each iteration is identical to the (inverse of) gradient of loss function $\mathcal{L}$. This observation is very useful since it implies that gradient EM is equivalent to gradient descent (GD) algorithm on $\mathcal{L}$. This observation is not a new discovery of ours but actually a wide-spread folklore (see [Jin et al., 2016]). However, our new contribution is to observe Fact 1 is very helpful for analyzing gradient EM, and to construct a new convergence analysis framework for gradient EM based on it.
1.4 Notation
In this paper, we adopt the following notational conventions. We denote $\{1,2,\ldots,n\}$ with $[n]$. $\bm{\mu}=(\mu_{1}^{\top},\ldots,\mu_{n}^{\top})^{\top}\in\mathbf{R}^{nd}$ denotes the parameter vector of GMM obtained by concatenating Gaussian mean vectors $\mu_{1},\ldots,\mu_{n}$ together. For any vector $\mu$, $\mu(t)$ denotes its value at time step $t$, sometimes we omit this iteration number $t$ when its choice is clear and simply abbreviate $\mu(t)$ as $\mu$. We define a shorthand of expectation taken over the ground truth GMM $\mathbf{E}_{x\sim\mathcal{N}(0,I_{d})}[\cdot]$ as $\mathbf{E}_{x}[\cdot]$. For any vector $v\neq 0$, we use $\overline{v}\coloneqq v/\|v\|$ to denote the normalization of $v$. We define (with a slight abuse of notation) ${i_{\max}}\coloneqq\arg\max_{i\in[n]}\{\|\mu_{i}\|\}$ as the index of $\mu_{i}$ with the maximum norm, and $\mu_{\max}\coloneqq\|\mu_{{i_{\max}}}\|=\max_{i\in[n]}\{\|\mu_{i}\|\}$ as the maximum norm of $\mu_{i}$. In particular, $\mu_{\max}(t)=\max\{\|\mu_{1}(t)\|,\ldots,\|\mu_{n}(t)\|\}$. Similarly, $\pi_{\min}\coloneqq\min_{i\in[n]}\pi_{i}$ and ${\pi_{\max}}\coloneqq\max_{i\in[n]}\pi_{i}$ denotes the minimal and maximal $\pi_{i}$, respectively. We use $\nabla_{\mu_{i}}\mathcal{L}$ to denote the gradient of $\mu_{i}$ on $\mathcal{L}$, and $\nabla\mathcal{L}=(\nabla_{\mu_{1}}\mathcal{L}^{\top},\ldots,\nabla_{\mu_{n}}\mathcal{L})^{\top}$ denotes the collection of all gradients. Finally we define a potential function $U:\mathbf{R}^{nd}\to\mathbf{R}\;$ for $\text{GMM}(\bm{\mu})$ as
$U(\bm{\mu})=\sum_{i\in[n]}\|\mu_{i}\|^{2}.$
@@ 1.5 Technical overview
New likelihood-based analysis framework. The traditional convergence analysis for EM/gradient EM in previous works Balakrishnan et al. [2014], Yan et al. [2017], Kwon and Caramanis [2020] proceeds by showing the distance between the model and the ground truth GMM in the parameter space contracts linearly in every iteration. This type of approach meets new challenges in the over-parameterized $n$-Gaussian mixture setting since the convergence is both sub-linear and non-monotonic. To address these problems, we propose a new likelihood-based convergence analysis framework: instead of proving the convergence of parameters, our analysis proceeds by showing the likelihood loss function $\mathcal{L}$ converges to 00. The new analysis framework is more flexible and allows us to overcome the aforementioned technical barriers.
@@ 2.2 N-Gaussian mixtures
A series of paper Yan et al. [2017], Zhao et al. [2018], Kwon and Caramanis [2020], Segol and Nadler follow the framework proposed by Balakrishnan et al. [2014] to prove the local convergence of EM for $n$-GMM. While their result applies to the more general $n$-Gaussian mixture ground truth setting, their framework only provides local convergence guarantee and cannot be directly applied to our setting.
@@ 3.1 Global convergence of gradient EM
Consider training a student $n$-component GMM initialized from $\bm{\mu}(0)=(\mu_{1}(0)^{\top},\ldots,\mu_{n}(0)^{\top})^{\top}$ to learn a single-component ground truth GMM $\mathcal{N}(0,I_{d})$ with population gradient EM algorithm. If the step size satisfies $\eta\leq O\left(\frac{\exp\left(-8U(0)\right)\pi_{\min}^{2}}{n^{2}d^{2}(\frac{1}{\mu_{\max}(0)}+\mu_{\max}(0))^{2}}\right)$, then gradient EM converges globally with rate
@@
where $\gamma=\Omega\left(\frac{\eta\exp\left(-16U(0)\right)\pi_{\min}^{4}}{n^{2}d^{2}(1+\mu_{\max}(0){\sqrt{dn}})^{4}}\right)\in\mathbf{R}^{+}$. Recall that $\mu_{\max}(0)=\max\{\|\mu_{1}(0)\|,\ldots,\|\mu_{n}(0)\|\}$ and $U(0)=\sum_{i\in[n]}\|\mu_{i}(0)\|^{2}$ are two initialization constants.
@@
The convergence rate in Theorem 2 has a factor exponentially small in the initialization scale ($\gamma\propto\exp(-16U(0))$). We would like to stress that this is again due to algorithmic nature of the problem rather than the limitation of analysis. In Section 3.2, we prove that there exists bad regions with exponentially small gradients so that when initialized from such region, gradient EM gets trapped locally for $\exp(\Omega(U(0)))$ number of steps. Therefore, a convergence speed guarantee exponentially small in $U(0)$ is inevitable and cannot be improved.
@@
Theorem 2 is fundamentally different from convergence analysis for EM/gradient EM in previous works Yan et al. [2017], Dwivedi et al. [2019], Balakrishnan et al. [2014] which proved monotonic linear contraction of parameter distance $\|\bm{\mu}(t)-\bm{\mu}^{*}\|$. But our result also implies global convergence since loss function $\mathcal{L}$ converging to 00 is equivalent to convergence of gradient EM to MLE.
@@ 3.2 Necessity of exponentially small factor in convergence rate
In this section we prove that a factor exponentially small in initialization scale ($\exp(-\Theta(U(0)))$) is inevitable in the global convergence rate guarantee of gradient EM. Particularly, we show the existence of bad regions such that initialization from this region traps gradient EM for exponentially long time before final convergence. Our result is the following theorem.
@@
For any $n\geq 3$, define $\tilde{\bm{\mu}}(0)=({\mu}^{\top}_{1}(0),\ldots,{\mu}^{\top}_{n}(0))$ as follows: ${\mu}_{1}(0)=12\sqrt{d}e_{1},{\mu}_{2}(0)=-12\sqrt{d}e_{1},\mu_{3}(0)=\cdots=\mu_{n}(0)=0$, where $e_{1}$ is a standard unit vector. Then population gradient EM initialized with means $\tilde{\bm{\mu}}(0)$ and equal weights $\pi_{1}=\ldots=\pi_{n}=1/n$ will be trapped in a bad local region around $\tilde{\bm{\mu}}(0)$ for exponentially long time
$T\coloneqq\frac{1}{30\eta}e^{d}=\frac{1}{30\eta}\exp(\Theta(U(0))).$
More rigorously, for any $0\leq t\leq T,\exists i\in[n]$ such that
$\|\mu_{i}(t)\|\geq 10\sqrt{d}.\;\;$
Theorem 7 states that, when initialized from some bad points $\bm{\mu}(0)$, after $\exp(\Theta(U(0)))$ number of time steps, gradient EM will still stay in this local region and remain $10\sqrt{d}$ distance away from the global minimum $\bm{\mu}=0$. Therefore an exponentially small factor in convergence rate is inevitable.
@@ 4.1 Difficulties of a global convergence proof and our new analysis framework
Proving the global convergence of gradient EM for general $n$-Gaussian mixture is highly nontrivial. While there have been many previous works [Balakrishnan et al., 2014, Yan et al., 2017, Dwivedi et al., 2018b] studying either local convergence or the special case of $2$-Gaussian mixtures, they all focus on showing the contraction of parametric error. Namely, their proof proceeds by showing the distance between the model parameter and the ground truth contracts, usually by a fixed linear ratio, in each iteration of the algorithm. However, this kind of approach faces various challenges for our general problem where the convergence is both sublinear and non-monotonic. Since the convergence rate is sublinear (see Remark 3), showing a linear contraction per iteration is no longer possible. Since the convergence is non-monotonic [^1] To see this, consider $n=2,\mu_{1}=0,\mu_{2}=(1,0,\ldots,0)^{\top}$, then the norm of $\mu_{1}$ strictly increases after one iteration., we also cannot show a strictly decreasing parametric distance.
@@ 4.2 Proof ideas for Theorem 2
Recall that $\bar{x}=\frac{x}{\|x\|}$. See detailed derivation in (23). Using (5), combined with the properties of $\nabla_{x}\tilde{\bm{\psi}}_{\bm{\mu}}(x)$, we can obtain the following lemma (Recall that $U=\sum_{i\in[n]}\|\mu_{i}\|^{2}$.):
@@
$\mathbf{E}_{x}\left[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}\right]\geq\frac{\exp\left(-8U\right)}{40000d(1+2\mu_{\max}{\sqrt{d}})^{2}}\left(\sum_{i,j\in[n]}\pi_{i}\pi_{j}\|\mu_{i}-\mu_{j}\|^{2}\right)^{2}.$
@@
$\left\langle\nabla_{\bm{\mu}}Q(\bm{\mu}|\bm{\mu}),\bm{\mu}\right\rangle=\mathbf{E}_{x}[\|\tilde{\bm{\psi}}_{\bm{\mu}}(x)\|^{2}]=\Omega\left(\frac{\exp\left(-8U\right)\pi_{\min}^{2}}{d(1+\mu_{\max}{\sqrt{d}})^{2}}\mu_{\max}^{4}\right).$
@@
The second problem is that our local smoothness theorem requires $\bm{\mu}$ to be bounded, therefore we need to show a regularity condition that for each $i$, $\bm{\mu}_{i}(t)$ stays in a bounded region during gradient EM updates. This is not easy to prove for each individual $\bm{\mu}_{i}$ due to the same non-monotonic issue mentioned in Section 4.1. To establish such a regularity condition, we use the potential function.$U$ to solve this problem. We prove that $U$ remains bounded along the gradient EM trajectory, implying each $\bm{\mu}_{i}$ remains well-behaved. With this regularity condition, combined with the previous two steps, we finish the proof of Theorem 2 via mathematical induction.